The ages of moe and joe now total 49 years. Moe's age now is twice the age that joe was when moe's age was the same as joe's is now. how old are moe and joe now?
I've tried doing this but I get mixed up when I get to the second sentence.
Thank you!

Answer by Edwin McCravy(20056)

The ages of moe and joe now total 49 years.
M+J = 49

Moe's age now is twice the age that joe was...
Stop here. Let y = the number of years ago that was.

y years ago, Moe was M-y and Joe was J-y

M = 2(J-y)

Now let's see what else was true y years ago:   

...when moe's age was the same as joe's is now.
M-y = J

So we have these three equations:

M+J = 49 
M = 2*(J-y)
M-y = J

Solve that system of equations and get J=21, M=28, y=7

Edwin